• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name B. Craver

    Log entry time 07:59:40 on March17,2005

    Entry number 141730

    keyword=Owl Shift Summary

    00:00 Begin shift during run 2732 with 40 uA on slanted carbon.
    05:30 MCC calls and reports they are having vacuum problems. No beam.
    05:53 MCC has solved the problem and the beam is back.
    07:52 Beam studies will begin at 8:30. I request that the hall be placed in Controlled Access at this time. After ~2.5 hrs MCC will need to continue studies with the hall in Beam Permit.
    08:00 End shift during run 2741

    Runs 2732 - 2737 slanted carbon (200k events)
    Run 2738 junk - beam is down
    Runs 2739,2740 slanted carbon 200k events
    Run 2741 slanted carbon - still running at end of shift